# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 3.3 Existence and uniqueness of weak solution
(3.12) $\|R_{1}(\widehat{p}_{1})-R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}\leq\dfrac{1}{c_{a}}\,\gamma\,\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}.$
Proof 3.9.
Let $(\boldsymbol{\sigma}_{1},(\boldsymbol{u}_{1},\boldsymbol{\rho}_{1})),\,(\boldsymbol{\sigma}_{2},(\boldsymbol{u}_{2},\boldsymbol{\rho}_{2}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))$, such that $\mathbf{R}(\widehat{p}_{1})=(\boldsymbol{\sigma}_{1},(\boldsymbol{u}_{1},\boldsymbol{\rho}_{1}))$ ad $\mathbf{R}(\widehat{p}_{2})=(\boldsymbol{\sigma}_{2},(\boldsymbol{u}_{2},\boldsymbol{\rho}_{2}))$. Then, from the definition of $\mathbf{R}$ (cf. (3.3)), we have
(3.13) $\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2},\boldsymbol{\tau})&+\quad\displaystyle b(\boldsymbol{\tau},(\boldsymbol{u}_{1}-\boldsymbol{u}_{2},\boldsymbol{\rho}_{1}-\boldsymbol{\rho}_{2}))&=\displaystyle-c(\boldsymbol{\tau},\widehat{p}_{1}-\widehat{p}_{2})&\forall\,\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2},(\boldsymbol{v},{\boldsymbol{\eta}}))&&=\displaystyle 0&\forall\,(\boldsymbol{v},{\boldsymbol{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega).\end{array}$
Since $\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2}\in\mathbb{V}$ (cf. (2.6a)), taking $\boldsymbol{\tau}=\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2}$ in (3.13), and utilising the ellipticity of $a$ on $\mathbb{V}$ (cf. (2.7)) along with the bound of $c$ (cf. (2.4b)), we obtain:
$c_{a}\|\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2}\|_{\mathbf{div};\Omega}^{2}\leq a(\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2},\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2})=-c(\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2},\widehat{p}_{1}-\widehat{p}_{2})\leq\,\gamma\,\|\boldsymbol{\sigma}_{1}-\boldsymbol{\sigma}_{2}\|_{\mathbf{div};\Omega}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega},$
which concludes the proof.
@@
Given $(\widehat{\boldsymbol{\sigma}}_{1},\,\widehat{p}_{1}),\,(\widehat{\boldsymbol{\sigma}}_{2},\,\widehat{p}_{2}),\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$, the following estimate holds
@@
(3.14) $\begin{array}[]{ll}\|S_{2}(\widehat{\boldsymbol{\sigma}}_{1},\,\widehat{p}_{1})-S_{2}(\widehat{\boldsymbol{\sigma}}_{2},\,\widehat{p}_{2})\|_{0,\Omega}\\[4.30554pt] \displaystyle\quad\leq\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}+\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}.\end{array}$
Proof 3.11.
@@
Let $({\boldsymbol{\varphi}}_{1},p_{1}),\,({\boldsymbol{\varphi}}_{2},p_{2})\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$, such that $\mathbf{S}(\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1})=({\boldsymbol{\varphi}}_{1},p_{1})$ and $\mathbf{S}(\widehat{\boldsymbol{\sigma}}_{2},\widehat{p}_{2})=({\boldsymbol{\varphi}}_{2},p_{2})$. Then, from the definition of $\mathbf{S}$ (cf. (3.4)), and employing similar arguments to those in Lemma 3.8, we have
$\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1}}({\boldsymbol{\varphi}}_{1},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})-\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{2},\widehat{p}_{2}}({\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})=-c(\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2},p_{1}-p_{2}),$
@@
by adding ±$\pm\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1}}({\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})$ in the last equation, we obtain
$\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1}}({\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})=\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{2},\widehat{p}_{2}}({\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})-\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1}}({\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})-c(\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2},p_{1}-p_{2}).$
@@
Then, using the the first assumption for $\kappa$ (cf. (2.3)), the ellipticity of $\widetilde{c}$ (see (2.9)), the definition of $\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}$ (cf. (2.1)) and the continuity of the form $c$ (see (2.4b)), we deduce
@@
$\begin{array}[]{ll}\displaystyle\kappa_{1}\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}^{2}+\widetilde{\gamma}\|p_{1}-p_{2}\|_{0,\Omega}^{2}\leq\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1}}({\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2},{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})\\[4.30554pt] \displaystyle\qquad=\int_{\Omega}(\kappa(\widehat{\boldsymbol{\sigma}}_{2},\widehat{p}_{2})^{-1}-\kappa(\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1})^{-1})\,{\boldsymbol{\varphi}}_{2}\cdot({\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2})-c(\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2},p_{1}-p_{2})\\[10.76385pt] \displaystyle\qquad\leq\|\kappa(\widehat{\boldsymbol{\sigma}}_{2},\widehat{p}_{2})^{-1}-\kappa(\widehat{\boldsymbol{\sigma}}_{1},\widehat{p}_{1})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\,\|{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\|p_{1}-p_{2}\|_{0,\Omega}.\end{array}$
@@
From the last equation, by utilising the second assumption regarding $\kappa$ (see (2.3)), we obtain
@@
$\begin{array}[]{ll}\frac{1}{2}\min\{\widetilde{\gamma},\kappa_{1}\}(\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega})^{2}\leq\kappa_{1}\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}^{2}+\widetilde{\gamma}\|p_{1}-p_{2}\|_{0,\Omega}^{2}\\[4.30554pt] \qquad\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\|p_{1}-p_{2}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad\leq\big{(}\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\big{(}\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega}\big{)},\end{array}$
@@
the last, together with the fact that ${\boldsymbol{\varphi}}_{2}$ satisfies (3.9), leads to the following bound
@@
$\begin{array}[]{ll}\frac{1}{2}\min\{\widetilde{\gamma},\kappa_{1}\}(\|{\boldsymbol{\varphi}}_{1}-{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega})\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\boldsymbol{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\\[4.30554pt] \displaystyle\qquad\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega})+\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega},\end{array}$
and this yields (3.14), concluding the proof.
@@
The following theorem presents the main result of this section, establishing the existence and uniqueness of the solution to the fixed-point problem (3.6), or equivalently, the well-posedness of problem (2.2).
@@
Given $r>0$, assume that $\boldsymbol{f}\in\mathbf{L}^{2}(\Omega)$, $g\in L^{2}(\Omega)$, $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1/2}(\Gamma_{\mathrm{D}})$, $p_{\mathrm{D}}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{D}})$ and $\gamma$ satisfies
(3.15) $\begin{array}[]{cc}\dfrac{2\mathop{\rm max}\limits\{1,\,\kappa_{2}\}}{\min\{\widetilde{\gamma},\,\kappa_{1}\,,r\}}\,\bigg{\{}\widetilde{C}(1+C_{1}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{\gamma^{2}}{c_{a}}(\dfrac{1}{\kappa_{2}}+\widetilde{C}\,r)\bigg{\}}<1.\end{array}$
Then, $\mathbf{T}$ (cf. (3.5)) has a unique fixed point $p\in\mathrm{W}$. Equivalently, (2.2) has a unique solution $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{W}$. In addition, there exists $C>0$, such that
(3.16) $\|\boldsymbol{\sigma}\|_{\mathbf{div};\Omega}+\|\boldsymbol{u}\|_{0,\Omega}+\|\boldsymbol{\rho}\|_{0,\Omega}+\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p\|_{0,\Omega}\leq C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,).$
Proof 3.13.
@@
$\leq\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}+\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|R_{1}(\widehat{p}_{1})-R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}$

### v2

@@ 3.3 Existence and uniqueness of weak solution
$\|R_{1}(\widehat{p}_{1})-R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}\leq\dfrac{1}{c_{a}}\,\gamma\,\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}.$ (3.12)
Proof.
Let $(\bm{\sigma}_{1},({\bm{u}}_{1},\bm{\rho}_{1})),\,(\bm{\sigma}_{2},({\bm{u}}_{2},\bm{\rho}_{2}))\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))$, such that $\mathbf{R}(\widehat{p}_{1})=(\bm{\sigma}_{1},({\bm{u}}_{1},\bm{\rho}_{1}))$ ad $\mathbf{R}(\widehat{p}_{2})=(\bm{\sigma}_{2},({\bm{u}}_{2},\bm{\rho}_{2}))$. Then, from the definition of $\mathbf{R}$ (cf. (3.3)), we have
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma}_{1}-\bm{\sigma}_{2},\bm{\tau})&+\displaystyle b(\bm{\tau},({\bm{u}}_{1}-{\bm{u}}_{2},\bm{\rho}_{1}-\bm{\rho}_{2}))&=\displaystyle-c(\bm{\tau},\widehat{p}_{1}-\widehat{p}_{2})&\forall\,\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\bm{\sigma}_{1}-\bm{\sigma}_{2},(\mathbf{v},{\bm{\eta}}))&&=\displaystyle 0&\forall\,(\mathbf{v},{\bm{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega).\end{array}$ (3.13)
Since $\bm{\sigma}_{1}-\bm{\sigma}_{2}\in\mathbb{V}$ (cf. (2.6a)), taking $\bm{\tau}=\bm{\sigma}_{1}-\bm{\sigma}_{2}$ in (3.13), and utilising the ellipticity of $a$ on $\mathbb{V}$ (cf. (2.7)) along with the bound of $c$ (cf. (2.4b)), we obtain:
$c_{a}\|\bm{\sigma}_{1}-\bm{\sigma}_{2}\|_{\mathbf{div};\Omega}^{2}\leq a(\bm{\sigma}_{1}-\bm{\sigma}_{2},\bm{\sigma}_{1}-\bm{\sigma}_{2})=-c(\bm{\sigma}_{1}-\bm{\sigma}_{2},\widehat{p}_{1}-\widehat{p}_{2})\leq\,\gamma\,\|\bm{\sigma}_{1}-\bm{\sigma}_{2}\|_{\mathbf{div};\Omega}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega},$
which concludes the proof. ∎
@@
Given $(\widehat{\bm{\sigma}}_{1},\,\widehat{p}_{1}),\,(\widehat{\bm{\sigma}}_{2},\,\widehat{p}_{2}),\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{W}$, the following estimate holds
@@
$\begin{array}[]{ll}\|S_{2}(\widehat{\bm{\sigma}}_{1},\,\widehat{p}_{1})-S_{2}(\widehat{\bm{\sigma}}_{2},\,\widehat{p}_{2})\|_{0,\Omega}\\[4.30554pt] \displaystyle\leq\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}+\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}.\end{array}$ (3.14)
Note that $\min\{\tilde{\gamma},\kappa_{1}\}$ enters in the denominator of the estimate, but we have assumed that $\kappa_{1}$ is strictly positive and it does not degenerate to zero.
Proof.
@@
Let $({\bm{\varphi}}_{1},p_{1}),\,({\bm{\varphi}}_{2},p_{2})\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$, such that $\mathbf{S}(\widehat{\bm{\sigma}}_{1},\widehat{p}_{1})=({\bm{\varphi}}_{1},p_{1})$ and $\mathbf{S}(\widehat{\bm{\sigma}}_{2},\widehat{p}_{2})=({\bm{\varphi}}_{2},p_{2})$. Then, from the definition of $\mathbf{S}$ (cf. (3.4)), and employing similar arguments to those in Lemma 3.6, we have
$\widetilde{a}_{\widehat{\bm{\sigma}}_{1},\widehat{p}_{1}}({\bm{\varphi}}_{1},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})-\widetilde{a}_{\widehat{\bm{\sigma}}_{2},\widehat{p}_{2}}({\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})=-c(\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2},p_{1}-p_{2}),$
@@
by adding ±$\pm\widetilde{a}_{\widehat{\bm{\sigma}}_{1},\widehat{p}_{1}}({\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})$ in the last equation, we obtain
$\widetilde{a}_{\widehat{\bm{\sigma}}_{1},\widehat{p}_{1}}({\bm{\varphi}}_{1}-{\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})$
$\qquad=\widetilde{a}_{\widehat{\bm{\sigma}}_{2},\widehat{p}_{2}}({\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})-\widetilde{a}_{\widehat{\bm{\sigma}}_{1},\widehat{p}_{1}}({\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})-c(\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2},p_{1}-p_{2}).$
@@
Then, using the the first assumption for ${\bm{\kappa}}$ (cf. (2.3)), the ellipticity of $\widetilde{c}$ (see (2.9)), the definition of $\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}$ (cf. (2.1)) and the continuity of the form $c$ (see (2.4b)), we deduce
@@
$\begin{array}[]{ll}\displaystyle\kappa_{1}\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}^{2}+\widetilde{\gamma}\|p_{1}-p_{2}\|_{0,\Omega}^{2}\leq\widetilde{a}_{\widehat{\bm{\sigma}}_{1},\widehat{p}_{1}}({\bm{\varphi}}_{1}-{\bm{\varphi}}_{2},{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})+\widetilde{c}(p_{1}-p_{2},p_{1}-p_{2})\\[4.30554pt] \displaystyle\qquad=\int_{\Omega}({\bm{\kappa}}(\widehat{\bm{\sigma}}_{2},\widehat{p}_{2})^{-1}-{\bm{\kappa}}(\widehat{\bm{\sigma}}_{1},\widehat{p}_{1})^{-1})\,{\bm{\varphi}}_{2}\cdot({\bm{\varphi}}_{1}-{\bm{\varphi}}_{2})-c(\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2},p_{1}-p_{2})\\[10.76385pt] \displaystyle\qquad\leq\|{\bm{\kappa}}(\widehat{\bm{\sigma}}_{2},\widehat{p}_{2})^{-1}-{\bm{\kappa}}(\widehat{\bm{\sigma}}_{1},\widehat{p}_{1})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\,\|{\bm{\varphi}}_{2}\|_{0,\Omega}\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\|p_{1}-p_{2}\|_{0,\Omega}.\end{array}$
@@
From the last equation, by utilising the second assumption regarding ${\bm{\kappa}}$ (see (2.3)), we obtain
@@
$\begin{array}[]{ll}\frac{1}{2}\min\{\widetilde{\gamma},\kappa_{1}\}(\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega})^{2}\leq\kappa_{1}\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}^{2}+\widetilde{\gamma}\|p_{1}-p_{2}\|_{0,\Omega}^{2}\\[4.30554pt] \qquad\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\bm{\varphi}}_{2}\|_{0,\Omega}\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\|p_{1}-p_{2}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad\leq\big{(}\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\bm{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\big{(}\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega}\big{)},\end{array}$
@@
the last, together with the fact that ${\bm{\varphi}}_{2}$ satisfies (3.9), leads to the following bound
@@
$\begin{array}[]{ll}\frac{1}{2}\min\{\widetilde{\gamma},\kappa_{1}\}(\|{\bm{\varphi}}_{1}-{\bm{\varphi}}_{2}\|_{0,\Omega}+\|p_{1}-p_{2}\|_{0,\Omega})\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\|{\bm{\varphi}}_{2}\|_{0,\Omega}+\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\\[4.30554pt] \displaystyle\qquad\leq\kappa_{2}\|\widehat{p}_{2}-\widehat{p}_{1}\|_{0,\Omega}\widetilde{C}\,(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega})+\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega},\end{array}$
and this yields (3.14), concluding the proof. ∎
@@
The following theorem presents the main result of this section, establishing the existence and uniqueness of the solution to the fixed-point problem (3.6), or equivalently, of (2.2).
@@
Given $r>0$, assume that $\bm{f}\in\mathbf{L}^{2}(\Omega)$, $g\in L^{2}(\Omega)$, ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1/2}(\Gamma_{\mathrm{D}})$, $p_{\mathrm{D}}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{D}})$ and $\gamma$ satisfies
$\begin{array}[]{cc}\dfrac{2\mathop{\rm max}\limits\{1,\,\kappa_{2}\}}{\min\{\widetilde{\gamma},\,\kappa_{1}\,,r\}}\,\bigg{\{}\widetilde{C}(1+C_{1}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{\gamma^{2}}{c_{a}}(\dfrac{1}{\kappa_{2}}+\widetilde{C}\,r)\bigg{\}}<1.\end{array}$ (3.15)
Then, $\mathbf{T}$ (cf. (3.5)) has a unique fixed point $p\in\mathrm{W}$. Equivalently, (2.2) has a unique solution $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{W}$. In addition, there exists $C>0$, such that
$\|\bm{\sigma}\|_{\mathbf{div};\Omega}+\|{\bm{u}}\|_{0,\Omega}+\|\bm{\rho}\|_{0,\Omega}+\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p\|_{0,\Omega}$
$\qquad\leq C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,).$ (3.16)
Proof.
@@
$\leq\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}$
$\qquad+\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|R_{1}(\widehat{p}_{1})-R_{1}(\widehat{p}_{2})\|_{\mathbf{div};\Omega}$